OpenStudy (anonymous):

HELP, please! Which expression is equivalent to (sin x + 1)(sin x − 1)? A. cos2x B. -cos2x C. cos2x + 1 D. cos2x − 1 E. -cos2x + 1

OpenStudy (anonymous):

\[=\sin^2x-1=-\cos^2x\]

OpenStudy (anonymous):

thank you can you help with one more please?

OpenStudy (anonymous):

yes,welcome.

OpenStudy (anonymous):

Simplify the expression: tan(-x)csc(x). Select one of the options below as your answer: A. cos x B. -cos x C. sin x D. sec x E. -sec x